### v1

@@ 4.4.1 Filtering
$\mathit{odt}_{\text{filtDep}}$ $\left(d,\mathit{xs}\right)=\mathit{min}_{E}\left(\mathit{filter}_{p}\left[\mathit{DN}\left(\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)$
@@
Proof.
The proof follows by substituting the fused generator and following the procedure of Theorems (3) and (4). For more general results and formal proofs, see He (2025). ∎
@@
The derivation for $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ similar to Theorem 6, but requires a different prefix-closed predicate definition, as the recursive structure of this two algorithms are sequential. S $e$ He (2025) for more general results.
@@ 4.4.2 Thinning algorithm (bounding technique)
Thinning is similar to, but distinct from, $\mathit{min}_{E}$. Indeed, the $\mathit{min}_{E}$ function can be understood as a special case of thinning with respect to a total order defined by the objective function $E$. In contrast, thinning is based on a preorder $R$, because transitivity is required, and total orders are too restrictive for the purposes of thinning (He, 2025; Bird and Gibbons, 2020). In a preorder relation, some configurations may be incomparable. Consequently, $\mathit{thin}_{R}:\left[\mathcal{A}\right]\to\left[\mathcal{A}\right]$ receives a list and returns a list, whereas $\mathit{min}_{E}:\left[\mathcal{A}\right]\to\mathcal{A}$ always returns a single element.
@@
For instance, in $\mathit{sodt}_{\text{vec}}$, the semantic interpretation of monotonicity in thinning is expressed as: $R\left(t,t^{\prime}\right)\Longrightarrow R\left(\mathit{update}\left(r,t\right),\mathit{update}\left(r,t^{\prime}\right)\right)$. Similarly, in $\mathit{sodt}_{\text{kperms}}$, we have: $R\left(\mathit{rs},\mathit{rs}^{\prime}\right)\Longrightarrow R\left(\mathit{insR}\left(r,\mathit{rs}\right),\mathit{insR}\left(r,\mathit{rs}^{\prime}\right)\right)$. Likewise, in $\mathit{sodt}_{\text{rec}}$, the semantic interpretation of monotonicity in thinning can be expressed as $R\left(u,u^{\prime}\right)\wedge R\left(v,v^{\prime}\right)\subseteq R\left(\mathit{DN}\left(u,r,v\right),\mathit{DN}\left(u^{\prime},r,v^{\prime}\right)\right)$.
@@
There are many ways to define a thinning algorithm He (2025), and its effectiveness depends on the context. For our purposes—the study of optimal decision tree problems—the dominance relation allows us to implement $\mathit{thin}$ simply using the $\mathit{filter}$ function. Therefore, we will not discuss the general construction of thinning, but instead focus on formalizing the specific dominance relations studied in ODT algorithms for machine learning research.
He (2025) showed that dominance relations commonly used in ML can be abstracted into two types: finite dominance relation (FDR) and global upper bound (GUB). Both relations rely on two auxiliary functions defined as follows.
Definition 5.
@@
Many existing studies on ODT problems focus on developing effective optimistic lower bounds or pessimistic upper bounds. He (2025) demonstrated that the following two generic dominance relations, derived using $\mathit{ub}$ and $\mathit{lb}$, satisfy the thinning laws. These abstract the dominance relations used in most optimal decision tree algorithms for machine learning.
Theorem 7.
@@
Proof.
To prove the dominance valid, we nee to verify the thinning algorithm defined using $\mathit{fdom}$ satisfies the universal property, which is equivalent to verifying the following monotonicity (Bird and De Moor, 1996):
@@
which holds because $\mathit{ub}^{*}\leq\mathit{lb}\left(a\right)\leq\mathit{lb}\left(f^{n}\left(a\right)\right)$. ∎
For the global upper bound dominance relation, the thinning function $\mathit{thin}:\left[\mathcal{A}\right]\to\left[\mathcal{A}\right]$ can be efficiently implemented using a simple $\mathit{filter}$ He (2025).
@@
We now introduce two useful and generic lower bounds proposed by Demirović et al. (2022) and Zhang et al. (2023) for classification and regression problems, respectively.
One particularly interesting and widely applicable bounding technique was introduced by Demirović et al. (2022); Lin et al. (2020); Brița et al. (2025) for classification problems. This technique is based on the following observation, although it relies on an assumption that has not yet been formally proven.
Definition 6.
@@
$E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right),\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)=E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|$
@@
$E_{\text{0-1}}\left(t_{\text{opt}}^{\prime},\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)<E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|,$
@@
$E_{2}\left(\mathit{sodt}\left(\mathit{xs},\mathit{rs}_{K}\right),\mathit{ys}\right)\geq\text{kmeans-1D}\left(K+1,\mathit{ys}\right)$
Proof.
The result follows directly, as a decision tree can be viewed as a method of clustering a one-dimensional data list—a label vector. Its clustering performance cannot surpass that of exact 1D $K$-means clustering, which is optimal with respect to squared loss. Consequently, any other clustering over $\mathit{ys}$, including that induced by the decision tree, will exhibit a greater squared loss. ∎
@@
Proof.
We prove 45 by following equational reasoning
@@
Note that the thinning algorithm does not necessarily guarantee a speedup compared to the original definition of $\mathit{odt}_{\text{depth}}$. Its performance depends on both the implementation of $\mathit{thin}$ and the effectiveness of the relation $R$. He (2025) provides various recursive definitions of $\mathit{thin}$ and the effectiveness of each definition is problem-dependent.
@@ 4.5.1 Comparison between size- and depth-constrained trees algorithms
As noted, depth-constrained decision trees are commonly used in ODT studies with relatively small combinatorics, such as axis-parallel trees (Mazumder et al., 2022; Brița et al., 2025) trees over binary-feature data (Demirović et al., 2022; Verwer and Zhang, 2019; Nijssen and Fromont, 2007; Aglin et al., 2020; Nijssen and Fromont, 2010). These existing methods typically process recursion in a depth-first manner, unlike the breadth-first formulation we defined in 34. In contrast, the size-constrained ODT algorithm has been far less studied; Hu et al. (2019); Lin et al. (2020); Zhang et al. (2023) have applied it to solve the ODT problem over binary-feature data, but the correctness of their algorithms remains unresolved.
@@
We first analyze the size-constrained decision tree and explain why it is suitable for parallel implementation. In 25, the original size-constrained ODT problem can be factorized as:
$\mathit{min}_{E}\circ\mathit{concatMapL}{}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}_{K}=\mathit{odt}_{K}$
This factorization indicates that solving the challenging $\mathit{odt}$ problem can be achieved by solving many smaller instances of $\mathit{sodt}$. As He and Little (2025a) demonstrated, there exists an elegant and efficient divide-and-conquer definition for $\mathit{kcombs}_{K}$, which decomposes the problem into smaller, independent subproblems. These subproblems can be processed in parallel and mapped efficiently onto modern memory hierarchies, particularly when combined with recursively blocked arrays (Gupta et al., 2000). Unfortunately, no similar factorization has yet been found for the depth-constrained ODT problem that would enable embarrassingly parallel execution. Consequently, the optimal decision tree algorithm derived from 23 can only be executed sequentially or parallelized post hoc, making it difficult to achieve the same level of speed-up as embarrassingly parallel algorithms.
@@
Although depth-first search effectively reduces memory usage, it significantly complicates parallelization compared to breadth-first strategies. Previous research on ODT problems with smaller combinatorics remains tractable even without parallelization because the number of possible splits at each recursive step is limited to $O\left(N\times D\right)$ for axis-parallel trees. However, this approach becomes ineffective when applied to the larger combinatorics of 23 where splitting rules are defined as hyperplanes. Hence, parallelization is likely the only viable solution for problems with such large combinatorial complexity.
@@ 5.1.1 Polynomial hypersurface and the $W$-tuple Veronese embedding
$P\left(x\right)=\sum_{i}w_{i}x^{\boldsymbol{\alpha}_{i}},w_{i}\in\mathbb{R}.$
@@
$\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)=2\sum_{d=0}^{D}\left(\begin{array}[]{c}N-1\\ d\end{array}\right).$
@@
$\mathit{minR}{}_{E_{\text{0-1}}}\left\{min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(D,\mathit{xs}\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{D}\right\}$
@@
Affine hyperplane $\left(x_{1},\ldots,x_{D},1\right)$ $w_{1}x_{1}+w_{2}x_{2}+\ldots+w_{D}x_{D}+1=0$, $D$
@@
Conic section $\left(x_{1},\ldots,x_{1}x_{D},\ldots,x_{D}^{2},1\right)$ $w_{1}x_{1}+\ldots w_{g}x_{1}x_{2}\ldots+w_{G}x_{D}^{2}+1=0$, $G=\left(\begin{array}[]{c}D+2\\ 2\end{array}\right)-1$
@@
Proof.
First, we need to prove the existence of the hypersurface. For any $G$ distinct points $\mathit{s}$ in $\mathbb{R}^{D+1}$, there exist a polynomial hypersurface $S$ defined by a polynomial $P$ with degree at most $M$ such that $\mathit{s}$ lies on $S$. This is because the vector space $V$ of a polynomial $P$ in $\mathbb{R}\left[x_{1},\ldots x_{D}\right]$of degree $M$ has a dimension $\left(\begin{array}[]{c}D+W\\ D\end{array}\right)$. For any data list $ds$ of size $G<\left(\begin{array}[]{c}D+M\\ D\end{array}\right)$, the linear map from $V$ to $\mathbb{R}^{G}$ must has non-trivial kernel (the linear system has more variables than equations), thus the polynomial hypersurface $S$ must exist. Next, consider a hyperplane $\boldsymbol{w}^{T}\bar{y}=0$ defined over variables $\bar{y}=\left(1,y_{1},\ldots y_{G}\right)\in\mathbb{R}^{G+1}$, we have $\boldsymbol{w}^{T}\bar{y}\geq$0 or $\boldsymbol{w}^{T}\bar{y}$¡0 implies $\boldsymbol{w}^{T}\rho_{M}\left(\bar{x}\right)\geq 0$ or $\boldsymbol{w}^{T}\rho_{M}\left(\bar{x}\right)<0$, because the closure of the real numbers under polynomial operation, thus $M\left(\bar{x}\right)\in\mathbb{R}$ if $\bar{x}\in\mathbb{R}^{D+1}$, for any monomial $M$. ∎
@@
Theorem 12.
0-1 loss polynomial hypersurface classification theorem. Consider a data list $\mathit{xs}$ of $N$ data points in $\mathbb{R}^{D}$ in general position, along with their associated labels. Denote $E_{\text{0-1}}\left(\boldsymbol{w}\right)$ as the 0-1 loss for evaluating normal vector $\boldsymbol{w}\in\mathbb{R}^{G}$ with respect to data lists $\rho_{M}\left(\mathit{xs}\right)$. Then the hyper degree $M$ polynomial
$\mathit{minR}{}_{E_{\text{0-1}}}\left\{min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(G,\rho_{M}\left(\mathit{xs}\right)\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{G}\right\}$

### v2

@@ 4.5.1 Filtering
$\quad\quad\left(\mathit{rs}^{+},r,\mathit{rs}^{-}\right)\leftarrow\mathit{splits}\left(\mathit{rs}\right)\big]\big).$
@@
$\mathit{odt}_{\text{filtDep}}$ $\left(d,\mathit{xs}\right)=\mathit{min}_{E}\big(\mathit{filter}_{p}\big[\mathit{DN}\left(\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid$
$\quad\quad r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\big]\big)$
@@
Proof The proof follows by substituting the fused generator and following the procedure of Theorems (7) and (8). For more general results and formal proofs, see He (2025b).
@@
The derivation for $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ similar to Theorem 14, but requires a different prefix-closed predicate definition, as the recursive structure of this two algorithms are sequential. See He (2025b) for more general results.
@@ 4.5.2 Thinning algorithm (bounding technique)
Thinning is similar to, but distinct from, $\mathit{min}_{E}$. Indeed, the $\mathit{min}_{E}$ function can be understood as a special case of thinning with respect to a total order defined by the objective function $E$. In contrast, thinning is based on a preorder $R$, because transitivity is required, and total orders are too restrictive for the purposes of thinning (He, 2025b; Bird and Gibbons, 2020). In a preorder relation, some configurations may be incomparable. Consequently, $\mathit{thin}_{R}:\left[\mathcal{A}\right]\to\left[\mathcal{A}\right]$ receives a list and returns a list, whereas $\mathit{min}_{E}:\left[\mathcal{A}\right]\to\mathcal{A}$ always returns a single element.
@@
For instance, in $\mathit{sodt}_{\text{vec}}$, the semantic interpretation of monotonicity in thinning is expressed as: $R\left(t,t^{\prime}\right)\Longrightarrow R\left(\mathit{update}\left(r,t\right),\mathit{update}\left(r,t^{\prime}\right)\right)$. Similarly, in $\mathit{sodt}_{\text{kperms}}$, we have: $R\left(\mathit{rs},\mathit{rs}^{\prime}\right)\Longrightarrow R\left(\mathit{insR}\left(r,\mathit{rs}\right),\mathit{insR}\left(r,\mathit{rs}^{\prime}\right)\right)$. Likewise, in $\mathit{sodt}_{\text{rec}}$, the semantic interpretation of monotonicity in thinning can be expressed as
$R\left(u,u^{\prime}\right)\wedge R\left(v,v^{\prime}\right)\subseteq R\left(\mathit{DN}\left(u,r,v\right),\mathit{DN}\left(u^{\prime},r,v^{\prime}\right)\right)$.
@@
There are many ways to define a thinning algorithm He (2025b), and its effectiveness depends on the context. For our purposes—the study of optimal decision tree problems—the dominance relation allows us to implement $\mathit{thin}$ simply using the $\mathit{filter}$ function. Therefore, we will not discuss the general construction of thinning, but instead focus on formalizing the specific dominance relations studied in ODT algorithms for machine learning research.
He (2025b) showed that dominance relations commonly used in ML can be abstracted into two types: finite dominance relation (FDR) and global upper bound (GUB). Both relations rely on two auxiliary functions defined as follows.
Definition 16
@@
Many existing studies on ODT problems focus on developing effective optimistic lower bounds or pessimistic upper bounds. He (2025b) demonstrated that the following two generic dominance relations, derived using $\mathit{ub}$ and $\mathit{lb}$, satisfy the thinning laws. These abstract the dominance relations used in most optimal decision tree algorithms for machine learning.
Theorem 17
@@
Proof To prove the dominance valid, we nee to verify the thinning algorithm defined using $\mathit{fdom}$ satisfies the universal property, which is equivalent to verifying the following monotonicity (Bird and De Moor, 1996):
@@
which holds because $\mathit{ub}^{*}\leq\mathit{lb}\left(a\right)\leq\mathit{lb}\left(f^{n}\left(a\right)\right)$.
For the global upper bound dominance relation, the thinning function $\mathit{thin}:\left[\mathcal{A}\right]\to\left[\mathcal{A}\right]$ can be efficiently implemented using a simple $\mathit{filter}$ He (2025b).
@@
We now introduce two useful and generic lower bounds proposed by Demirovic et al. (2022) and Zhang et al. (2023) for classification and regression problems, respectively.
One particularly interesting and widely applicable bounding technique was introduced by Demirovic et al. (2022); Lin et al. (2020); Brita et al. (2025) for classification problems. This technique is based on the following observation, although it relies on an assumption that has not yet been formally proven.
Definition 18
@@
$E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right),\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)=E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|$ (54)
@@
$E_{\text{0-1}}\left(t_{\text{opt}}^{\prime},\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)<E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|,$ (55)
@@
$E_{2}\left(\mathit{sodt}\left(\mathit{xs},\mathit{rs}_{K}\right),\mathit{ys}\right)\geq\text{kmeans-1D}\left(K+1,\mathit{ys}\right)$ (56)
Proof The result follows directly, as a decision tree can be viewed as a method of clustering a one-dimensional data list—a label vector. Its clustering performance cannot surpass that of exact 1D $K$-means clustering, which is optimal with respect to squared loss. Consequently, any other clustering over $\mathit{ys}$, including that induced by the decision tree, will exhibit a greater squared loss.
@@
$\mathit{odt}_{\text{thindep}}$ $\left(1,\mathit{xs}\right)=\mathit{min}_{E}\left[\mathit{DN}\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right]$
@@
Proof We prove 57 by following equational reasoning
@@
Note that the thinning algorithm does not necessarily guarantee a speedup compared to the original definition of $\mathit{odt}_{\text{depth}}$. Its performance depends on both the implementation of $\mathit{thin}$ and the effectiveness of the relation $R$. He (2025b) provides various recursive definitions of $\mathit{thin}$ and the effectiveness of each definition is problem-dependent.
@@ 4.6.1 Comparison between size- and depth-constrained trees algorithms
As noted, depth-constrained decision trees are commonly used in ODT studies with relatively small combinatorics, such as axis-parallel trees (Mazumder et al., 2022; Brita et al., 2025) trees over binary-feature data (Demirovic et al., 2022; Verwer and Zhang, 2019; Nijssen and Fromont, 2007; Aglin et al., 2020; Nijssen and Fromont, 2010). These existing methods typically process recursion in a depth-first manner, unlike the breadth-first formulation we defined in 37. In contrast, the size-constrained ODT algorithm has been far less studied; Hu et al. (2019); Lin et al. (2020); Zhang et al. (2023) have applied it to solve the ODT problem over binary-feature data, but the correctness of their algorithms remains unresolved.
@@
We first analyze the size-constrained decision tree and explain why it is suitable for parallel implementation. In 26, the original size-constrained ODT problem can be factorized as:
$\mathit{min}_{E}\circ\mathit{concatMapL}{}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}_{K}=\mathit{odt}_{K}$ (58)
This factorization indicates that solving the challenging $\mathit{odt}$ problem can be achieved by solving many smaller instances of $\mathit{sodt}$. As He and Little (2025a) demonstrated, there exists an elegant and efficient divide-and-conquer definition for $\mathit{kcombs}_{K}$, which decomposes the problem into smaller, independent subproblems. These subproblems can be processed in parallel and mapped efficiently onto modern memory hierarchies, particularly when combined with recursively blocked arrays (Gupta et al., 2000). Unfortunately, no similar factorization has yet been found for the depth-constrained ODT problem that would enable embarrassingly parallel execution. Consequently, the optimal decision tree algorithm derived from 24 can only be executed sequentially or parallelized post hoc, making it difficult to achieve the same level of speed-up as embarrassingly parallel algorithms.
@@
Although depth-first search effectively reduces memory usage, it significantly complicates parallelization compared to breadth-first strategies. Previous research on ODT problems with smaller combinatorics remains tractable even without parallelization because the number of possible splits at each recursive step is limited to $O\left(N\times D\right)$ for axis-parallel trees. However, this approach becomes ineffective when applied to the larger combinatorics of 24 where splitting rules are defined as hyperplanes. Hence, parallelization is likely the only viable solution for problems with such large combinatorial complexity.
@@ 5.1.1 Polynomial hypersurface and the $W$-tuple Veronese embedding
$P\left(x\right)=\sum_{i}w_{i}x^{\boldsymbol{\alpha}_{i}},w_{i}\in\mathbb{R}.$ (59)
@@
$\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)=2\sum_{d=0}^{D}\left(\begin{array}[]{c}N-1\\ d\end{array}\right).$ (60)
@@
$\mathit{minR}_{E_{\text{0-1}}}\left\{\min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(D,\mathit{xs}\right)\right\}\subseteq$ (62)
$\quad\mathit{minR}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)\right\}\subseteq\mathit{minR}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{D}\right\}$
@@
Affine hyperplane $\left(x_{1},\ldots,x_{D},1\right)$ $w_{1}x_{1}+w_{2}x_{2}+\ldots+w_{D}x_{D}+1=0$$w\bar{\boldsymbol{x}}=0$, $D$
@@
Conic section $\left(x_{1},\ldots,x_{1}x_{D},\ldots,x_{D}^{2},1\right)$ $w_{1}x_{1}+\ldots w_{g}x_{1}x_{2}\ldots+w_{G}x_{D}^{2}+1=0$$w\bar{\boldsymbol{x}}=0$, $G=\left(\begin{array}[]{c}D+2\\ 2\end{array}\right)-1$
@@
Proof First, we need to prove the existence of the hypersurface. For any $G$ distinct points $\mathit{s}$ in $\mathbb{R}^{D+1}$, there exist a polynomial hypersurface $S$ defined by a polynomial $P$ with degree at most $M$ such that $\mathit{s}$ lies on $S$. This is because the vector space $V$ of a polynomial $P$ in $\mathbb{R}\left[x_{1},\ldots x_{D}\right]$of degree $M$ has a dimension $\left(\begin{array}[]{c}D+W\\ D\end{array}\right)$. For any data list $xs$ of size $G<\left(\begin{array}[]{c}D+M\\ D\end{array}\right)$, the linear map from $V$ to $\mathbb{R}^{G}$ must has non-trivial kernel (the linear system has more variables than equations), thus the polynomial hypersurface $S$ must exist. Next, consider a hyperplane $\boldsymbol{w}^{T}\bar{y}=0$ defined over variables $\bar{y}=\left(1,y_{1},\ldots y_{G}\right)\in\mathbb{R}^{G+1}$, we have $\boldsymbol{w}^{T}\bar{y}\geq$0 or $\boldsymbol{w}^{T}\bar{y}$¡0 implies $\boldsymbol{w}^{T}\rho_{M}\left(\bar{x}\right)\geq 0$ or $\boldsymbol{w}^{T}\rho_{M}\left(\bar{x}\right)<0$, because the closure of the real numbers under polynomial operation, thus $M\left(\bar{x}\right)\in\mathbb{R}$ if $\bar{x}\in\mathbb{R}^{D+1}$, for any monomial $M$.
@@
Theorem 25
0-1 loss polynomial hypersurface classification theorem. Consider a data list $\mathit{xs}$ of $N$ data points in $\mathbb{R}^{D}$ in general position, along with their associated labels. Denote $E_{\text{0-1}}\left(\boldsymbol{w}\right)$ as the 0-1 loss for evaluating normal vector $\boldsymbol{w}\in\mathbb{R}^{G}$ with respect to data lists $\rho_{M}\left(\mathit{xs}\right)$. Then we have
$\mathit{minR}{}_{E_{\text{0-1}}}\left\{min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(G,\rho_{M}\left(\mathit{xs}\right)\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{G}\right\}$ (63)